40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 1 16:01:48 2011, Jamie, Update, SUS, ETM binary whitening switching fixed 
    Reply  Fri Jul 1 18:54:34 2011, Jamie, Update, SUS, ETMY binary whitening switching confirmed to be fixed 
Message ID: 4932     Entry time: Fri Jul 1 18:54:34 2011     In reply to: 4929
Author: Jamie 
Type: Update 
Category: SUS 
Subject: ETMY binary whitening switching confirmed to be fixed 

After finally figuring out what was messed up with ETMY I was able to get good measurements of the binary whitening switching on ETMY to determine that it is in fact working now:

ETMY
ul : 3.2937569959  = 10.3538310999 db
ll : 3.28988426634 = 10.3436124066 db
sd : 3.34670033732 = 10.4923365497 db
lr : 3.08727050163 =  9.7914936665 db
ur : 3.27587751842 = 10.3065531117 db

ELOG V3.1.3-